# The numeraire e-variable

### 29 Feb 2024 v1

@@ 3 Duality and reverse information projection
$H({\mathbb{Q}}\mid{\mathbb{P}})={\mathbb{E}}_{\mathbb{Q}}\left[-\log\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right].$ (3.1)
@@
Remark 3.3.
In the above, ${\mathbb{P}}$ can be any element of $M_{+}$. We do not adjust the definition of relative entropy for the possibility that its total mass is not one.
We now make the trivial observation that
${\mathbb{E}}_{\mathbb{Q}}\left[X\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right]\leq 1\text{ for all $X\in{\mathcal{E}}$ and ${\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$.}$ (3.2)
Indeed, the left-hand side equals ${\mathbb{E}}_{{\mathbb{P}}^{a}}[X]$ which is dominated by ${\mathbb{E}}_{\mathbb{P}}[X]$, hence by one. This leads to the following theorem, which expresses the duality between the null hypothesis and its set of e-variables and justifies calling ${\mathbb{P}}^{*}$ the RIPr. We first treat the notationally simpler and statistically more relevant case where ${\mathbb{Q}}\ll{\mathcal{P}}$ and, hence, the numeraire $X^{*}$ is finite. The general case is in Theorem 3.7.
@@
Assume ${\mathbb{Q}}\ll{\mathcal{P}}$ and let $X^{*}$ be the numeraire e-variable and ${\mathbb{P}}^{*}$ the RIPr. Writing as above ${\mathbb{P}}^{a}$ for the absolutely continuous part of ${\mathbb{P}}$ with respect to ${\mathbb{Q}}$ we have
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\frac{d{\mathbb{P}}^{a}}{d{\mathbb{P}}^{*}}\right]\leq 1\text{ for all }{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ},$ (3.3)
and ${\mathbb{P}}^{*}$ is the unique element of ${\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathbb{Q}}$ that satisfies (3.3). Furthermore, (3.3) is equivalent to
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\log\frac{d{\mathbb{P}}^{a}}{d{\mathbb{P}}^{*}}\right]\leq 0\text{ for all }{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}.$ (3.4)
Finally, one has the strong duality relation
${\mathbb{E}}_{\mathbb{Q}}[\log X^{*}]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathbb{Q}}[\log X]=\inf_{{\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}}H({\mathbb{Q}}\mid{\mathbb{P}})=H({\mathbb{Q}}\mid{\mathbb{P}}^{*}),$ (3.5)
where some, and then all, of these quantities may be +$\infty$+ ∞. Here ${\mathbb{E}}_{\mathbb{Q}}[\log X]$ is understood as $\infty$- ∞ whenever ${\mathbb{E}}_{\mathbb{Q}}[(\log X)^{-}]=\infty$.
The property (3.3) satisfied by ${\mathbb{P}}^{*}$ can be viewed as dual to the numeraire property and generalizes Property 2 in Li (1999, Theorem 4.3). The equivalence of (3.3) and (3.4) is analogous to the equivalence between the numeraire property and log-optimality of the numeraire established in Proposition 2.4. Finally, (3.5) states that strong duality holds between the problem of maximizing the expected logarithm over e-variables and the problem of minimizing relative entropy over the effective null. Note that ${\mathbb{E}}_{\mathbb{Q}}[\log X^{*}]$ is well-defined and nonnegative, possibly infinite, thanks to the numeraire property and the fact that $\log X^{*}\geq-1/X^{*}+1$. Note also that (3.5) implies that there exists an e-variable with positive expected logarithm under ${\mathbb{Q}}$ if and only if ${\mathbb{Q}}\notin{\mathcal{P}}^{\circ\circ}$. This complements a result by Zhang et al. (2023), which shows that if ${\mathcal{P}}$ is a polytope (the convex hull of a finite number of distributions) and there exists a common reference measure, then a such an e-variable exists if and only if ${\mathbb{Q}}$ is not in the convex hull of ${\mathcal{P}}$.
The minimum relative entropy property of ${\mathbb{P}}^{*}$ justifies calling it the RIPr of ${\mathbb{Q}}$ on ${\mathcal{P}}^{\circ\circ}$ as in Definition 3.2. In the terminology of Lardy et al. (2023), (3.4) implies that the description gain of switching from ${\mathbb{P}}^{*}$ to any other element of ${\mathcal{P}}^{\circ\circ}$ is nonpositive, so that ${\mathbb{P}}^{*}$ is their so-called universal reverse information projection of ${\mathbb{Q}}$ on ${\mathcal{P}}^{\circ\circ}$. Existing literature on the (universal) RIPr assumes the existence of a reference measure along with much stronger versions of the condition that ${\mathbb{Q}}\ll{\mathcal{P}}$. As we show in Section 4, in such settings our ${\mathbb{P}}^{*}$ coincides with existing notions of the (universal) RIPr of ${\mathbb{Q}}$ on ${\mathcal{P}}$; see Theorem 4.6. However, our measure ${\mathbb{P}}^{*}$ exists in complete generality and satisfies properties normally associated with the RIPr; see Theorem 3.7 below for the fully general case. This leads us to Definition 3.2, which we argue is the natural definition of the RIPr in the absence of any assumptions.
Proof of Theorem 3.4.
Since ${\mathbb{P}}^{*}$ and ${\mathbb{Q}}$ are equivalent under the given assumption, we have $d{\mathbb{P}}^{a}/d{\mathbb{P}}^{*}=(d{\mathbb{P}}^{a}/d{\mathbb{Q}})/(d{\mathbb{P}}^{*}/d{\mathbb{Q}})=X^{*}d{\mathbb{P}}^{a}/d{\mathbb{Q}}$ up to ${\mathbb{Q}}$-nullsets for any ${\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$. We thus obtain (3.3) by taking $X=X^{*}$ in (3.2). Next, the uniqueness statement is argued exactly as uniqueness of the numeraire: if ${\mathbb{P}}_{1}^{*},{\mathbb{P}}_{2}^{*}\in{\mathcal{P}}^{\circ\circ}$ are both equivalent to ${\mathbb{Q}}$ and satisfy (3.3), then with $Y=d{\mathbb{P}}^{*}_{2}/d{\mathbb{P}}^{*}_{1}$ we have $1\leq 1/{\mathbb{E}}_{\mathbb{Q}}[Y]\leq{\mathbb{E}}_{\mathbb{Q}}[1/Y]\leq 1$ so that $Y$ is ${\mathbb{Q}}$-almost surely equal to one and hence ${\mathbb{P}}^{*}_{1}={\mathbb{P}}^{*}_{2}$. Further, (3.3) implies (3.4) thanks to Jensen’s inequality, and the converse uses the same argument as the proof of Proposition 2.4. Specifically, ${\mathbb{P}}(t)=(1-t){\mathbb{P}}^{*}+t{\mathbb{P}}$ belongs to ${\mathcal{P}}^{\circ\circ}$ for all $t\in(0,1)$, so ${\mathbb{E}}_{\mathbb{Q}}[t^{-1}\log(d{\mathbb{P}}(t)^{a}/d{\mathbb{P}}^{*})]\leq 0$. Sending $t$ to zero yields (3.3). Finally, we argue (3.5). Jensen’s inequality and (3.2) give
@@
${\mathbb{E}}_{\mathbb{Q}}\left[\log\left(X\frac{d{\mathbb{P}}^{a}}{d{\mathbb{Q}}}\right)\right]\leq 0$ (3.6)
for all $X\in{\mathcal{E}}$ and ${\mathbb{P}}\in{\mathcal{P}}^{\circ\circ}$. This implies the ‘weak duality’ inequality

### 22 Mar 2024 v2

@@ 3 Duality and reverse information projection
${\mathbb{E}}_{\mathsf{Q}}\left[X\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right]\leq 1\text{ for all $X\in{\mathcal{E}}$ and ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$.}$ (3.2)
Indeed, the left-hand side equals ${\mathbb{E}}_{{\mathsf{P}}^{a}}[X]$ which is dominated by ${\mathbb{E}}_{\mathsf{P}}[X]$, hence by one. This leads to the following theorem, which expresses the duality between the null hypothesis and its set of e-variables and justifies calling ${\mathsf{P}}^{*}$ the RIPr. We first treat the notationally simpler and statistically more relevant case where ${\mathsf{Q}}\ll{\mathcal{P}}$ and, hence, the numeraire $X^{*}$ is finite. The general case is in Theorem 3.8.
Theorem 3.4.
Assume ${\mathsf{Q}}\ll{\mathcal{P}}$, let $X^{*}$ be the numeraire e-variable, and let ${\mathsf{P}}^{*}$ be an element of ${\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathsf{Q}}$. The following statements are equivalent, where as above we write ${\mathsf{P}}^{a}$ for the absolutely continuous part of ${\mathsf{P}}$ with respect to ${\mathsf{Q}}$ (and hence ${\mathsf{P}}^{*}$).
  1. (i)
    ${\mathsf{P}}^{*}$ is the RIPr.
  2. (ii)
    ${\mathbb{E}}_{\mathsf{Q}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$.
  3. (iii)
    ${\mathbb{E}}_{\mathsf{Q}}\left[\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}.$
If any of these hold, one has the strong duality relation
${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[\log X]=\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}H({\mathsf{Q}}\mid{\mathsf{P}})=H({\mathsf{Q}}\mid{\mathsf{P}}^{*}),$ (3.3)
where some, and then all, of these quantities may be +$\infty$+ ∞. Here ${\mathbb{E}}_{\mathsf{Q}}[\log X]$ is understood as $\infty$- ∞ whenever ${\mathbb{E}}_{\mathsf{Q}}[(\log X)^{-}]=\infty$.
Property (ii) satisfied by the RIPr can be viewed as dual to the numeraire property and generalizes Property 2 in Li (1999, Theorem 4.3). The equivalence of (ii) and (iii) is analogous to the equivalence between the numeraire property and log-optimality of the numeraire established in Proposition 2.4. Finally, (3.3) states that strong duality holds between the problem of maximizing the expected logarithm over e-variables and the problem of minimizing relative entropy over the effective null. Note that ${\mathbb{E}}_{\mathsf{Q}}[\log X^{*}]$ is well-defined and nonnegative, possibly infinite, thanks to the numeraire property and the fact that $\log X^{*}\geq-1/X^{*}+1$.
The minimum relative entropy property of ${\mathsf{P}}^{*}$ justifies calling it the RIPr of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$ as in Definition 3.2. In the terminology of Lardy et al. (2023), property (ii) implies that the description gain of switching from ${\mathsf{P}}^{*}$ to any other element of ${\mathcal{P}}^{\circ\circ}$ is nonpositive, so that ${\mathsf{P}}^{*}$ is their universal reverse information projection of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$. Existing literature on the (universal) RIPr assumes the existence of a reference measure along with much stronger versions of the condition that ${\mathsf{Q}}\ll{\mathcal{P}}$. As we show in Section 4, in such settings our ${\mathsf{P}}^{*}$ coincides with existing notions of the (universal) RIPr of ${\mathsf{Q}}$ on ${\mathcal{P}}$; see Theorem 4.7 and its corollaries. However, our measure ${\mathsf{P}}^{*}$ exists in complete generality and satisfies properties normally associated with the RIPr; see Theorem 3.8 below for the fully general case. This leads us to Definition 3.2, which we argue is the natural definition of the RIPr in the absence of any assumptions.
Proof of Theorem 3.4.
(i) $\Leftrightarrow$ (ii): Suppose ${\mathsf{P}}^{*}$ is the RIPr. Because ${\mathsf{P}}^{*}$ and ${\mathsf{Q}}$ are equivalent we have $d{\mathsf{P}}^{a}/d{\mathsf{P}}^{*}=(d{\mathsf{P}}^{a}/d{\mathsf{Q}})/(d{\mathsf{P}}^{*}/d{\mathsf{Q}})=X^{*}d{\mathsf{P}}^{a}/d{\mathsf{Q}}$ up to ${\mathsf{Q}}$-nullsets for any ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$. Thus we obtain (ii) by taking $X=X^{*}$ in (3.2). This proves the forward direction. To prove the converse, we show that the RIPr is the unique element of ${\mathcal{P}}^{\circ\circ}$ that is equivalent to ${\mathsf{Q}}$ and satisfies (ii). This is argued exactly as uniqueness of the numeraire: if ${\mathsf{P}}_{1}^{*},{\mathsf{P}}_{2}^{*}\in{\mathcal{P}}^{\circ\circ}$ are both equivalent to ${\mathsf{Q}}$ and satisfy ${\mathbb{E}}_{\mathsf{Q}}[{d{\mathsf{P}}^{a}}/{d{\mathsf{P}}_{i}^{*}}]\leq 1$ for all ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$, then with $Y=d{\mathsf{P}}^{*}_{2}/d{\mathsf{P}}^{*}_{1}$ we have $1\leq 1/{\mathbb{E}}_{\mathsf{Q}}[Y]\leq{\mathbb{E}}_{\mathsf{Q}}[1/Y]\leq 1$. Thus $Y$ is ${\mathsf{Q}}$-almost surely equal to one and ${\mathsf{P}}^{*}_{1}={\mathsf{P}}^{*}_{2}$.
(ii) $\Leftrightarrow$ (iii): The forward implication follows from Jensen’s inequality, and the converse uses the same argument as the proof of Proposition 2.4. Specifically, ${\mathsf{P}}(t)=(1-t){\mathsf{P}}^{*}+t{\mathsf{P}}$ is in ${\mathcal{P}}^{\circ\circ}$ for all $t\in(0,1)$, so ${\mathbb{E}}_{\mathsf{Q}}[t^{-1}\log(d{\mathsf{P}}(t)^{a}/d{\mathsf{P}}^{*})]\leq 0$. Sending $t$ to zero yields (ii).
Finally, we argue (3.3) assuming that ${\mathsf{P}}^{*}$ is the RIPr. Jensen’s inequality and (3.2) give
${\mathbb{E}}_{\mathsf{Q}}\left[\log\left(X\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)\right]\leq 0$ (3.4)
for all $X\in{\mathcal{E}}$ and ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$. This implies the ‘weak duality’ inequality
${\mathbb{E}}_{\mathsf{Q}}\left[\log X\right]\leq{\mathbb{E}}_{\mathsf{Q}}\left[-\log\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right]=H({\mathsf{Q}}\mid{\mathsf{P}}).$ (3.5)
Indeed, if $(\log X)^{-}$ has infinite expectation, (3.5) is trivially true. Otherwise we first replace $X$ by $X\wedge n$, in which case (3.4) can be rearranged to (3.5) with $X\wedge n$ in place of $X$. Sending $n$ to infinity and using monotone convergence then gives (3.5). Since equality is achieved by $X=X^{*}$ and ${\mathsf{P}}={\mathsf{P}}^{*}$, we deduce (3.3). ∎
Remark 3.5.
The strong duality in (3.3) implies that there exists an e-variable with positive expected logarithm under ${\mathsf{Q}}$ if and only if ${\mathsf{Q}}\notin{\mathcal{P}}^{\circ\circ}$. Therefore, thanks to Corollary 4.6 below, Theorem 3.4 extends a result by Zhang et al. (2023) showing that if ${\mathcal{P}}$ is finite, then such an e-variable exists if and only if ${\mathsf{Q}}$ is not in the convex hull of ${\mathcal{P}}$.
Remark 3.6.
The RIPr ${\mathsf{P}}^{*}$ is not a probability measure in general. Equivalently, it may happen that ${\mathbb{E}}_{\mathsf{Q}}[1/X^{*}]<1$. We will see this phenomenon in Sections 5.1 and 5.4, and other examples are given by Lardy et al. (2023). It is also worth noting that there may exist elements $\widetilde{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$ that are different from ${\mathsf{P}}^{*}$, but whose absolutely continuous part $\widetilde{\mathsf{P}}^{a}$ is equal to ${\mathsf{P}}^{*}$. We will see this in Section 5.4 with $\widetilde{\mathsf{P}}$ being the symmetrization of ${\mathsf{Q}}$. Other examples are obtained from Example 3.7 below.
Example 3.7 (Example 2.2 continued).
Consider a simple null hypothesis ${\mathcal{P}}=\{{\mathsf{P}}_{0}\}$ with ${\mathsf{Q}}\ll{\mathsf{P}}_{0}$ so that the numeraire $X^{*}$ is finite. In this case ${\mathsf{P}}^{*}={\mathsf{P}}_{0}^{a}$ and $X^{*}=1/(d{\mathsf{P}}^{*}/d{\mathsf{Q}})=d{\mathsf{Q}}/d{\mathsf{P}}_{0}$ up to ${\mathsf{Q}}$-nullsets. Here one can easily check that ${\mathcal{P}}^{\circ\circ}$ consists of all sub-probability measures ${\mathsf{P}}$ such that ${\mathsf{P}}(A)\leq{\mathsf{P}}_{0}(A)$ for all $A$. Indeed, any such ${\mathsf{P}}$ must belong to ${\mathcal{P}}^{\circ\circ}$ since ${\mathbb{E}}_{\mathsf{P}}[X]\leq{\mathbb{E}}_{{\mathsf{P}}_{0}}[X]\leq 1$ for all e-variables $X$. Conversely, if ${\mathsf{P}}$ is not of this form, then ${\mathsf{P}}(A)>{\mathsf{P}}_{0}(A)$ for some $A$, and $X=\bm{1}_{A}/{\mathsf{P}}_{0}(A)$ is an e-variable such that ${\mathbb{E}}_{\mathsf{P}}[X]>1$, showing that ${\mathsf{P}}\notin{\mathcal{P}}^{\circ\circ}$. If ${\mathsf{P}}_{0}(A)=0$ we set $X=\infty\bm{1}_{A}$ in the previous sentence.
We end this section with the general version of Theorem 3.4. The proof reduces the general case to the one treated in Theorem 3.4 by introducing the conditional probability measure ${\mathsf{Q}}^{*}(A)={\mathsf{Q}}(A\mid X^{*}<\infty)$ if the quantity $\lambda^{*}={\mathsf{Q}}(X^{*}<\infty)$ is positive, where $X^{*}$ is the numeraire. The fully degenerate case $\lambda^{*}=0$ is treated separately. It is evident from Definition 3.2 that the RIPr ${\mathsf{P}}^{*}$ is absolutely continuous with respect to ${\mathsf{Q}}$, and we know from Lemma 3.1 that it belongs to ${\mathcal{P}}^{\circ\circ}$. As shown below, one actually has more precise information, namely that
@@
where we write $\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$ for the set $\{\lambda^{*}{\mathsf{P}}\colon{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}\}\subset{\mathcal{P}}^{\circ\circ}$. We continue to write ${\mathsf{P}}^{a}$ for the absolutely continuous part of ${\mathsf{P}}$ with respect to ${\mathsf{Q}}$.
@@
Let $X^{*}$ be the numeraire e-variable and define $\lambda^{*}={\mathsf{Q}}(X^{*}<\infty)$. The RIPr ${\mathsf{P}}^{*}$ satisfies (3.6). Furthermore, for any sub-probability measure ${\mathsf{P}}^{*}$ that is absolutely continuous with respect to ${\mathsf{Q}}$ and satisfies (3.6), the following statements are equivalent.
  1. (i)
    ${\mathsf{P}}^{*}$ is the RIPr.
  2. (ii)
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\bm{1}_{\{X^{*}<\infty\}}\right]\leq 1\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}.$
  3. (iii)
@@
    ${\mathbb{E}}_{\mathsf{Q}}\left[\log\left(\lambda^{*}\frac{d{\mathsf{P}}^{a}}{d{\mathsf{P}}^{*}}\right)\bm{1}_{\{X^{*}<\infty\}}\right]\leq 0\text{ for all }{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}.$
If any of these hold, one has the strong duality relation (3.3). Finally, suppose $\lambda^{*}\in(0,1)$, let ${\mathsf{P}}^{*}$ be the RIPr, and define ${\mathsf{Q}}^{*}={\mathsf{Q}}(\cdot\mid X^{*}<\infty)$. Then the numeraire and RIPr associated with ${\mathsf{Q}}^{*}$ are $X^{*}$ and ${\mathsf{P}}^{**}={\mathsf{P}}^{*}/\lambda^{*}$, and one has the ‘conditional’ strong duality
${\mathbb{E}}_{{\mathsf{Q}}^{*}}[\log X^{*}]=\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{{\mathsf{Q}}^{*}}[\log X]=\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}H({\mathsf{Q}}^{*}\mid{\mathsf{P}})=H({\mathsf{Q}}^{*}\mid{\mathsf{P}}^{**}).$ (3.7)
Proof.
We first let ${\mathsf{P}}^{*}$ be the RIPr and check (3.6). Equivalence with ${\mathsf{Q}}$ on $\{X^{*}<\infty\}$ is evident from Definition 3.2. To show that ${\mathsf{P}}^{*}$ belongs to $\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$ we refine the calculation in the proof of Lemma 3.1. Consider any $X\in{\mathcal{E}}$ and set $X^{\prime}=X\bm{1}_{\{X^{*}<\infty\}}+\infty\bm{1}_{\{X^{*}=\infty\}}$. This still belongs to ${\mathcal{E}}$ since it is ${\mathsf{P}}$-almost surely equal to $X$ for all ${\mathsf{P}}\in{\mathcal{P}}$. Thanks to the convention (1.1) we have $X^{\prime}/X^{*}=\bm{1}_{\{X^{*}<\infty\}}X^{\prime}/X^{*}+\bm{1}_{\{X^{*}=\infty\}}$ and hence
${\mathbb{E}}_{{\mathsf{P}}^{*}}[X]={\mathbb{E}}_{{\mathsf{P}}^{*}}[X^{\prime}]={\mathbb{E}}_{\mathsf{Q}}\left[\bm{1}_{\{X^{*}<\infty\}}\frac{X^{\prime}}{X^{*}}\right]={\mathbb{E}}_{\mathsf{Q}}\left[\frac{X^{\prime}}{X^{*}}\right]-{\mathsf{Q}}(X^{*}=\infty)\leq{\mathsf{Q}}(X^{*}<\infty),$
using the numeraire property in the last step. It follows that ${\mathsf{P}}^{*}\in\lambda^{*}\cdot{\mathcal{P}}^{\circ\circ}$.
We now let ${\mathsf{P}}^{*}$ be any sub-probability measure that is absolutely continuous with respect to ${\mathsf{Q}}$ and satisfies (3.6), and prove the claimed equivalences and the strong duality statement. First, in the fully degenerate case $\lambda^{*}=0$, (3.6) just states that ${\mathsf{P}}^{*}=0$. This is also the RIPr, so (i) is true. Moreover, (ii) and (iii) reduce to the statements that $0\leq 1$ and $0\leq 0$. Thus all three statements are true and and hence trivially equivalent. The strong duality (3.3) holds because all four terms are infinite. Second, in the ‘fully non-degenerate’ case $\lambda^{*}=1$, the theorem simply reduces to Theorem 3.4, which we have already proved.
Consider now the case $\lambda^{*}\in(0,1)$ and define ${\mathsf{Q}}^{*}={\mathsf{Q}}(\cdot\mid X^{*}<\infty)$. Taking ${\mathsf{Q}}^{*}$ as our new alternative hypothesis, $X^{*}$ is still the numeraire. Indeed, it is a ${\mathsf{Q}}^{*}$-almost surely strictly positive e-variable, and for any other e-variable $X$ we define $X^{\prime}$ as above and get
@@
${\mathbb{E}}_{{\mathsf{Q}}^{*}}\left[\frac{X}{X^{*}}\right]=\frac{1}{\lambda^{*}}{\mathbb{E}}_{\mathsf{Q}}\left[\bm{1}_{\{X^{*}<\infty\}}\frac{X}{X^{*}}\right]=\frac{1}{\lambda^{*}}\left({\mathbb{E}}_{\mathsf{Q}}\left[\frac{X^{\prime}}{X^{*}}\right]-{\mathsf{Q}}(X^{*}=\infty)\right)\leq 1.$
Since ${\mathsf{Q}}^{*}(X^{*}<\infty)=1$ by construction, we have from Theorem 2.5 that ${\mathsf{Q}}^{*}\ll{\mathcal{P}}$. We also define ${\mathsf{P}}^{**}={\mathsf{P}}^{*}/\lambda^{*}$ which, thanks to (3.6), is an element of ${\mathcal{P}}^{\circ\circ}$ equivalent to ${\mathsf{Q}}^{*}$. Now, the statement (i) that ${\mathsf{P}}^{*}$ is the RIPr of ${\mathsf{Q}}$ is equivalent to ${\mathsf{P}}^{**}$ being the RIPr of ${\mathsf{Q}}^{*}$. Furthermore, (ii) and (iii) can be equivalently expressed in terms of ${\mathsf{Q}}^{*}$ and ${\mathsf{P}}^{**}$ as